If x-4=11 then the value of X-8 is?
Found 2 solutions by macston, fractalier:
Answer by macston(5194)

.
x-4=11
x=15
.
x-8=15-8=7
.
ANSWER: x-8=7
.

Answer by fractalier(6550)

If we know that
x-4=11
then we can see how
x = 15
so that
x - 8 = 15 - 8 = 7